OpenStudy (anonymous):

what are the graphing points of y=2x+1

jimthompson5910 (jim_thompson5910):

let's say we plug in x = 0

jimthompson5910 (jim_thompson5910):

y=2x+1 y=2(0)+1 ... replace every x with 0 (since x = 0) y = 0 + 1 y = 1 When x = 0, the value of y is y = 1 They pair up to form the point (0,1)

jimthompson5910 (jim_thompson5910):

What is the value of y when x = 1 ?